Cesar Rego Computers, a British Columbia chain of computer hardware and software retail outlets, supplies both educational and commercial customers with memory and storage devices. It currently faces the following ordering decision relating to purchases of high-density disks:

D = 36 000 disks

S = $25

H = $0.45

Purchase price = $0.85

Discount price = $0.82

Quantity needed to qualify for the discount = 6000 disks Should the discount be taken
Answer #1

As per the given question Annual Demand (D)=36000 disks

cost per order(S)=$25

Holding Cost per disk (H)=$0.45

Economic Order Quantity= √2DS/H

=√2*36000*25/0.45

=√180000/0.45

=√4000000

=2000

EOQ minimise the inventory cost

We need to compare the inventory cost at various levals in case of discount

so we have to caluculate inventory cost for 2000 and 6000 disks

EOQ after applying discount =√2*36000*25/0.82+0.45

=√180000/1.27

=√1417323

=1190.51

=1190 disks

Order Quantity 1190 disks 2000 Disks 6000 Disks
No of orders (Annual demand/Order Quantity) 36000/1190 =30.25 36000/2000=18 36000/6000=6
Ordering cost (No of orders*$25) 30.25*$25=$756.25 18*$25=$450 6*25= $150
Holding Cost ($0.45*Average Inventory) $0.45*1190/2 = $267.75 $0.45*2000/2=$450 $0.45*6000/2=$1350
Cost of Purchase (Annual demand*Discount Price) 36000*$0.82 = $29520 36000*$0.82 = $29520 36000*0.82= $29520
Total Inventory Cost 30544 30420 31020

Based on the above analysis optimum order leval is 2000 disks